Shift Summary for owl shift, 01/13/2000


Subject: Shift Summary for owl shift, 01/13/2000
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:22:11 CST


              Shift Summary for owl shift, 01/13/2000

On Shift: J. Shields, A. Tripathi

Hours of beam available at all intensities: 7.0
Hours of beam available at nominal intensity: 7.0
Hours of E799 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 6.4

Accounting of lost time:

starting/stopping runs, swapping tapes
========================================================================
Runs ended this shift:

Run 15498 -- SHORT_REG832
Run 15499 -- SHORT_REG832
Run 15500 -- SHORT_REG832

Tapes brought to FCC this shift:
  (including tapes for online split)

QKE101-109, QKB159
========================================================================
Accesses during shift:

none

========================================================================
Overview of shift:

Relatively smooth short regenerator running. We had some DAQ_DEAD and
L3 accept fraction problems after we stopped run 15498 to swap tapes,
but 2 brief runs later we recovered.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

15498 00:00 inherit run from prev. shift. Spill shape is a
                        bit erratic, but apparently much better than
                        it was earlier in the evening.

              00:30 MCR calls and promises to tune up the spill
                        shape.
 
              00:35 spill shape looks good, as promised.

              05:00 Sydney says "boy I'm full." Stopped run, swapped
                        tapes, and did a drt_fresh_start. Mounted
                        QKE110-118, QKB160.

15499 05:12 begin new run. Changed #hit requirement for
                        Trigs 11,12 from 33 to 40.

              05:27 DAQ_DEAD alarms and L3 accept fraction probs
                        on ktev2 plane. Crew decides to stop run and
                        change Trig 11,12 requirement back to 33.

15500 05:34 start new run; forced L3 namelist

              05:39 L3 probs still refuse to go away. Crew rolls
                        their eyeballs in annoyance and decide to stop
                        this run as well and then do another
                        drt_fresh_start

15501 06:00 Start yet ANOTHER new run. Third time apparently
                        really *IS* a charm, since L3 problems have
                        finally gone away. More smooth running ensues.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:22:13 CST